OpenStudy (hhopke):

WILL MEDAL AND FAN! What are the horizontal and vertical intercepts of the following functions? ~f(x)=3√x-3 *cube root of (x-3) ~f(x)=3√29x+15 ~g(x)=4√-12x+16 *fourth root of (-12x+16) ~f(x)=5√-x *fifth root of (-x) ~p(x)=6√-24-27x *sixth root of (-24-27x)

OpenStudy (hhopke):

The following answers aren't correct- ~x=3, y=-1.442 ~x=-0.517, y=2.466 ~x=1.333, y=2 ~x=0, y=0 ~x=8/-9 *the correct answer for y is (y=DNE)

OpenStudy (anonymous):

i have never heard of a 'horizontal intercept" before, but maybe that means where it crosses the \(x\) axis does that sound right?

OpenStudy (hhopke):

yup! :)

OpenStudy (anonymous):

and a vertical one is where it crosses the \(y\) axis then i guess

OpenStudy (anonymous):

the vertical one will be easiest to find, since on the \(y\) axis you have \(x=0\) so to find it for each one, compute \(f(0)\)

OpenStudy (anonymous):

for example \[f(x)=\sqrt[3]{x-3}\] then \[f(0)=\sqrt[3]{0-3}=\sqrt[3]{-3}\] if you need a decimal use a calculator

OpenStudy (anonymous):

for the horiizontal one, set \(f(x)=0\) and solve for \(x\) for example if \[f(x)=\sqrt[3]{x-3}\] then set \[\sqrt[3]{x-3}=0\] cube both sides to get \[x-3=0\] and so \[x=3\]

OpenStudy (hhopke):

The only problem is, with the first one, I put \[\sqrt[3]{-3}\] into my calculator and it gave me -1.442, which is the answer that I got the first time and was marked wrong. :(

OpenStudy (hhopke):

Same thing with the x-intercept one.

OpenStudy (anonymous):

then i don't know what to say, because they look right to me maybe they want the answer written as an ordered pair can you post a screen shot?

OpenStudy (hhopke):

The attach file thing is buggy on my computer. You're probably right about the ordered pair thing, though. :)

OpenStudy (hhopke):

It worked! Got the file thing to work. Thanks so much for your help! :) <3

OpenStudy (anonymous):

yay! yw